A train X leaves,Meerut at 5:00am and reaches Delhi at 9:00am.Another trainY
leaves Delhi at 7:00am and reaches Meerut at 10:30am.At what time do two
trains cross each others

Answer by lwsshak3(11628)

A train X leaves,Meerut at 5:00am and reaches Delhi at 9:00am.Another trainY
leaves Delhi at 7:00am and reaches Meerut at 10:30am.At what time do two
trains cross each others
***
let d=distance between Meerut and Delhi
speed of train x=distance/ total travel time=d/4
speed of train y=distance/total travel time=d/3.5
let x=travel time of train y when the 2 trains cross each other
x+2=travel time of train x when the 2 trains cross each other
(x+2)(d/4)+x(d/3.5)=d (sum of distances traveled by each train=total distance between Meerut and Delhi)
d's cancel out
(x+2)/4+x/3.5=1
3.5(x+2)+4x=4*3.5=14
3.5x+7+4x=14
7.5x=7
x=7/7.5≈0.93 hr=56 min
x+2≈2.93=2 hrs and 56 min
Check:
2.93/4≈73% of total distance between Meerut and Delhi
0.93/3.5=27% of total distance between Meerut and Delhi
At what time do two trains cross each other? 7:56 am